P=12N, Q=5N, theta=120 degrees
Find the resultant force and alpha
1
Expert's answer
2020-02-27T10:13:07-0500

As per the given question,

P=12N

Q=5N

Angle between the force "\theta=120^\circ"

"R=\sqrt{P^2+Q^2+2PQ\cos\theta}"


"R=\sqrt{12^2+5^2+2\times12\times5\cos(120^\circ)}"

"R=\sqrt{144+25+120\cos{(90^\circ+30^\circ)}}"

"R=\sqrt{169-120\sin30^\circ}"

"R=\sqrt{169-120\dfrac{1}{2}}" N

"R=\sqrt{169-60}=\sqrt{109}"